A boat travels 100 miles downstream in the same length of time it travels 50 miles upstream. If the boat's speed and still water at 30 mph, find the speed of the current of the river.
Answer by mananth(16946)

Boat speed = 30 mph
current speed = x mph

speed against current 30 - x
speed with current 30 + x

Distance against current 50 miles
Distance with current 100 miles

t=d/r time against = time with current -
50 / ( 30 - x )= 100 / ( 30 + x )
50 ( 30 + x )= 100 ( 30 -x )
1500 + 50 x = 3000 - -100 x
100 x 50 x 3000 + -1500
150 x = 1500
/ 150
x = 10 mph current speed